(5 pts) Suppose I sell 5 PUT option contracts on Home Depot with a strike price...

  1. (5 pts) Suppose I sell 5 PUT option contracts on Home Depot with a strike price of $180.00 and an option premium of $19. The stock price is $209.45 when I sell the put and moves between $199.24 and $215.43 during the time up to expiration. What is my maximum profit or loss on this?
  1. (5 pts) Using the date in question 13, what would my maximum profit or loss be if the price of Home Depot stock moved from $172 to $225.45?

  1. (5 pts) The option premium for a call on Apple with a strike price of $240 is $5.60 for a December 2019 expiration, but the premium is $4.20 for a call with the same strike price and a November 2019 expiration date. Why is the option premium less for the November expiration?
Answer #1

1.
Profit at 215.43=-5*100*(MAX(180-215.43,0)-19)=9500
Profit at 199.24=-5*100*(MAX(180-199.24,0)-19)=9500

Hence, maximum profit is 9500

2.
Profit at 172=-5*100*(MAX(180-172,0)-19)=5500
Profit at 225.45=-5*100*(MAX(180-225.45,0)-19)=9500

Hence, maximum profit is 9500

3.
Option premium increases with maturity